Transcript Lecture 14
UNR, MATH/STAT 352, Spring 2007 Binomial(n,p) n k n! n k P( X k ) p (1 p ) p k (1 p ) n k (n k )! k ! k E ( X ) np np(1 p ) 2 X UNR, MATH/STAT 352, Spring 2007 0.1 0.2 0.3 Number of successes within 1 symmetric Bernoulli trial can only be 0 or 1. These possibilities have equal chances. 0.0 dbinom(q, 1, 0.5) 0.4 0.5 Binomial(1,0.5) 0 2 4 6 Number of succsses UNR, MATH/STAT 352, Spring 2007 8 10 0.5 Binomial(2,0.5) 0.3 0.2 0.1 dbinom(q, 2, 0.5) 0.4 Number of successes within 2 symmetric Bernoulli trials can only be 0, 1 or 2. The possibility to have exactly 1 success is larger than that of having 0 or 2. 0.0 A fair game should result in a tie. 0 2 4 6 8 10 Then why do people play fair games? Number of succsses UNR, MATH/STAT 352, Spring 2007 0.8 Binomial(2,0.1) 0.4 0.2 0.0 dbinom(q, 2, 0.1) 0.6 Most likely there will be no successes 0 2 4 6 Number of succsses UNR, MATH/STAT 352, Spring 2007 8 10 0.8 Binomial(2,0.9) 0.4 0.2 0.0 dbinom(q, 2, 0.9) 0.6 Most likely there will be only successes 0 2 4 6 Number of succsses UNR, MATH/STAT 352, Spring 2007 8 10 0.20 0.25 Unimodal (mode = 1) Right-skewed (E = np = 1.5) 0.05 0.10 0.15 Concentrated around 1.5 0.00 dbinom(q, 15, 0.1) 0.30 0.35 Binomial(15,0.1) 0 2 4 6 Number of succsses UNR, MATH/STAT 352, Spring 2007 8 10 Unimodal (mode = 1-2) 0.20 Right-skewed 0.05 0.10 0.15 Concentrated around 1.5 0.00 dbinom(q, 19, 0.1) 0.25 Binomial(19,0.1) 0 2 4 6 Number of succsses UNR, MATH/STAT 352, Spring 2007 8 10 0.12 Unimodal (mode = 10) 0.10 Symmetric? (E = np = 10) 0.08 Concentrated around 10 0.02 0.04 0.06 P(10+3) < P(10-3) 0.00 dbinom(q, 100, 0.1) Binomial(100,0.1) 0 5 10 15 20 Number of succsses UNR, MATH/STAT 352, Spring 2007 25 30 0.12 0.10 Symmetric? (E = np = 90) Concentrated around 90 0.02 0.04 0.06 0.08 Unimodal (mode = 90) 0.00 dbinom(q, 100, 0.9) Binomial(100,0.9) 70 75 80 85 90 Number of succsses UNR, MATH/STAT 352, Spring 2007 95 100 Bin(100,0.9) 0.04 0.06 0.08 0.10 Only a small fraction of possible outcomes has not negligible P (i.e. only small part can be seen in experiment) 0.02 P is very small (not 0!) here 0.00 dbinom(q, 100, 0.1) 0.12 Bin(100,0.1) 0 20 40 60 Number of succsses UNR, MATH/STAT 352, Spring 2007 80 100 0.30 Binomial(6,0.5) 0.20 0.15 0.10 0.05 0.00 dbinom(q, 6, 0.5) 0.25 Here all possible outcomes have reasonable probabilities 0 1 2 3 4 Number of succsses UNR, MATH/STAT 352, Spring 2007 5 6 Poisson(l) P( X k ) e l E( X ) l lk k! l 2 X UNR, MATH/STAT 352, Spring 2007 Poisson(1) 0.3 0.2 0.1 0.0 0.2 0 2 4 6 8 0.1 Number of succsses 0.0 dpois(q, 1) 0.3 dbinom(q, 1000, 1/1000) Binomial(1000,.001] 0 2 4 6 Number of succsses 8 I’ve seen this already! UNR, MATH/STAT 352, Spring 2007 10 10 n + p 0 + np l Binomial(n,p) Poisson(l) UNR, MATH/STAT 352, Spring 2007 Poisson(30) Binomial(1000,.03) N(30,sqrt(30)) 0.06 Poisson 0.02 0.04 Normal 0.00 dpois(q, 30) Binomial 0 10 20 30 Number of succsses UNR, MATH/STAT 352, Spring 2007 40 50 If n is large (n > 100), p is small (p < 0.05), and both np and n(1-p) are not small (say >10) then B(n,p)~P(np)~N(np, np(1-p)) UNR, MATH/STAT 352, Spring 2007 0.7 Normal densities 0.5 0.6 Mean= 0, Std=0.6 0.4 0.3 Mean= 0, Std=1 0.1 0.2 Mean= 5, Std=1 Mean= 0, Std=3 0.0 Density Mean= 7, Std=1 -10 -8 -6 -4 -2 0 2 4 Value of random variable UNR, MATH/STAT 352, Spring 2007 6 8 10 0.006 Interval probability Lower-tail probability Upper-tail probability 0.000 Probability 0.012 Probability density function 2400 2450 2500 2550 2600 Value of random variable 0.8 0.4 0.0 Probability Cumulative distribution function 2400 2450 2500 2550 Value of random variable UNR, MATH/STAT 352, Spring 2007 2600 UNR, MATH/STAT 352, Spring 2007 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 2 games. P(lose or win < $10)= 1 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 4 games. P(lose or win < $10)= 1 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 6 games. P(lose or win < $10)= 1 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 10 games. P(lose or win < $10)= 1 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 100 games. P(lose or win < $10)= 0.728 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 1000 games. P(lose or win < $10)= 0.272 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100 Payoff 2 X N , where X Bin(N , 1 2 ) (Why this formula? See lecture notes.) 0.3 0.2 0.1 0.0 Probability of payoff 0.4 0.5 10000 games. P(lose or win < $10)= 0.087 -100 -50 0 50 Net payoff, in $ UNR, MATH/STAT 352, Spring 2007 100